# Detecting Masquerade Attacks in Controller Area Networks Using Graph Machine Learning

### v1

@@ IV-B Time Series Extraction and Analysis
Concurrently with the creation of the MSGs, raw CAN messages are decoded using a DBC as described in [63]. In doing so, CAN data is transformed into a time series format to capture detailed patterns of the signals contained in every CAN ID per sliding window. We used CAN-D [64] to extract timeseries from CAN logs. Up to the time of this writing, CAN-D is still the state-of-the-art method for CAN reverse engineering [65]. Fig. 5 shows a decoded representation of the signals representing the four wheels’ speed of a vehicle [15]. The plot is generated by translating the CAN frames into a continuous time series format, which captures the network activity over time. This representation facilitates a detailed characterization of masquerade attacks within CANs highlighting variations and trends that may not be evident relying on raw CAN data.
Figure 5: Time series encoded in node ID 1760, illustrating signal patterns of four wheels’ speed. In this figure, the x-axis represents time and the y-axis represents the values of signals through time.
@@ IV-C Graph Annotation
We enrich MSGs further by annotating each node with statistical attributes derived from the time series data. In particular, we include the mean and standard deviation of each signal associated with each CAN ID. Note that other time series features can be computed but we opted for simpler statistical features that have constant complexity [3]. Our hypothesis is that enriching MSGs with contextual data from the time series significantly enhances our ability to conduct graph ML and detect subsequent masquerade attacks. Note that the graph annotation process is performed as MSGs become available. This approach ensures that CAN communications’ temporal and semantic aspects are captured comprehensively. Fig. 6 shows how each node is annotated with statistical attributes, specifically the mean ($\mu$) and standard deviation ($\sigma$) of the associated signals. Algorithm 2 describes the pseudocode of this process.
@@
Algorithm 2 Node annotation with statistical attributes
0:  Data from CAN, list of sliding windows, list of signals
0:  Graph with nodes annotated with statistical attributes
1:  for each sliding window in the list of sliding windows do
2:     Define the start and end of the current sliding window
3:     for each signal in the list of signals do
4:        Extract signal values within the current sliding window
5:        Interpolate missing values using linear interpolation
6:        Calculate the $\mu$ and $\sigma$ of the interpolated signal
7:        Assign calculated $\mu$ and $\sigma$ to the corresponding node in the graph
8:     end for
9:  end for
10:  return  Annotated graph
@@ IV-D Graph Embeddings
Graph embeddings provide a powerful way to represent the complex structure of graphs into a lower dimensional space such that ML algorithms can effectively take advantage [66, 67, 68]. In our study, we focus on shallow graph embeddings and in particular the node2vec algorithm [39], which is well-suited for capturing the nuances of weighted and directed graphs characteristic of MSG. Moreover, node2vec facilitates the transformation of MSG nodes into vector representations, preserving nodes’network topology. We configure node2vec with specific parameters—64 dimensions (d), a walk length (l) of 15, 100 walks (r) per node, a return parameter (p) of 1.5, and an in-out parameter (q) of 0.5. This configuration allows capturing the unique characteristics of MSGs while excelling at inference tasks included in the graph annotation phase [39]. Our configuration ensures that the embedding process remains efficient, thereby supporting near real-time detection capabilities, which is critical for timely identification of masquerade attacks in CANs. Note that this relatively low embedding dimension has shown to be enough for capturing rich graph-related properties [69]. In our work, instead of focusing on node embeddings, we construct whole-graph embeddings [32]. We compute them by averaging node embeddings within sliding windows leading to a unified representation of the network’s behavior. Specifically, we stack them per node in the MSGs. Each of these graphs is labeled indicating the normal or anomalous activity in CAN, i.e., we frame our IDS problem as a graph classification problem [70]. Algorithm 3 describes how we generate our node2vec embeddings.
Algorithm 3 node2vec embedding generation
0:  List of graphs for each time slice, node2vec parameters (d, l, r, p, q)
0:  List of node embeddings with associated statistical attributes for each graph
1:  for each graph in the list of time slice graphs do
2:     Initialize node2vec with the specified parameters (d, l, r, p, q)
3:     Fit node2vec model to the graph
4:     for each node in the graph do
5:        Retrieve or initialize the embedding vector for the node
6:        Annotate the embedding vector with statistical attributes
7:     end for
8:     Average the embedding vectors to form a consistent network representation
9:  end for
10:  return  DataFrame of averaged node embeddings with labels
@@ IV-E Supervised Learning
To evaluate the effectiveness of our feature representation and overall anomaly detection framework, we employ supervised learning techniques. We experiment with two ensemble learning methods: Random Forest (RF) [71] and XGBoost [72]. These decision tree-based ensemble learning methods are renowned for their high accuracy, scalability, and robustness, making them exceptionally suitable for dealing with the high-dimensional space characteristic of CAN data. The RF model is instantiated with 100 trees, and depth is capped at 20 to prevent overfitting. A combination of a minimum samples split of 6 and minimum samples leaf of 2 is used to control the growth of the trees further. As the training dataset exhibits a 1:10 ratio between the minority (attack) and majority (normal) classes, we balance the dataset using the synthetic minority over-sampling technique (SMOTE) [73]. A 5-fold stratified k-fold cross-validation is performed to ensure the model’s generalizability during training. The XGBoost model yielded similar results to the RF classifier in our experiments. Despite the comparable performance, we opted to report only results for the RF classifier due to its simpler interpretability and less intensive hyperparameter tuning process, which aligns well with the goals of this study.
@@ IV-F Evaluation Setup
Evaluation is performed at the sliding window level. In doing so, we verify the proportion of sliding windows coinciding with an attack region in the testing captures, which refers to the recorded CAN bus data used for testing. We found that for the combinations of $\omega$ and $\delta$, the proportion of windows containing attacks tended to be balanced, i.e., greater than 40%. To quantitatively assess the performance of our anomaly detection framework, we use the area under the receiver operating characteristic curve (AUC-ROC) [74]. The AUC-ROC is a comprehensive evaluation metric for classification problems across various threshold settings. It reflects the model’s ability to distinguish between classes, with higher values indicating better performance. The ROC curve is a plot of the true positive rate (TPR) against the false positive rate (FPR) at different threshold levels. The TPR, also known as sensitivity or recall, measures the proportion of actual positives correctly identified, while the FPR measures the proportion of actual negatives incorrectly identified as positive. Mathematically, the AUC-ROC is defined as:
[…]
The value of the AUC-ROC ranges between 0 and 1. An AUC-ROC of 1 indicates perfect classification, whereas an AUC-ROC of 0.5 suggests no discriminative power, equivalent to random guessing. An AUC-ROC below 0.5 indicates that the model performs worse than random guessing. In practical terms, the AUC-ROC can be interpreted as the probability that the classifier will rank a randomly chosen positive instance higher than a randomly chosen negative instance. This interpretation is particularly useful in our context, where we aim to detect subtle masquerade attacks embedded within normal CAN traffic.
@@ IV-G Dataset
Our unifying framework is evaluated on the Real ORNL Automotive Dynamometer (ROAD) dataset, a comprehensive set of CAN data collected from a real vehicle at Oak Ridge National Laboratory (ORNL) [15]. This dataset is particularly valuable due to the inclusion of physically verified fabrication and simulated masquerade attacks that provide a realistic environment for testing CAN security methods. The ROAD dataset stands out for its unparalleled quality of CAN information covering the most credible CAN attacks reflecting various driving scenarios. The dataset comprises 3.5 hours of recorded data, with 3 hours allocated for training and the remaining 30 minutes set aside for testing. The test data incorporates five masquerade attacks including ‘correlated signal’, ‘max engine’, ‘max speedometer’, ‘reverse light off’, and ‘reverse light on’ attacks. These attacks are designed to manipulate specific vehicle states. This dataset enables us to conduct a comprehensive evaluation of our framework under a variety of attack conditions. Table I provides a comprehensive overview of the masquerade attacks used in our study and their impacts on the vehicle’s state.
TABLE I: Characteristics of masquerade attacks in the ROAD dataset.
Attack name Attack details and consequences
Correlated signal Injects varying values for wheel speeds, resulting in the vehicle coming to a halt.
Max speedometer Injects the maximum value to be displayed on the speedometer.
Max coolant temp Injects the maximum value, triggering the coolant warning light.
Reverse light on/off Toggles the reverse light irrespective of the actual gear position.
@@ V Results
Our experimental setup was powered by a 12th Gen Intel® Core i9-12900HK processor, an NVIDIA GeForce RTX 3080 Ti GPU, and bolstered by 32GB of RAM. We employed NetworkX, a powerful library for creating and analyzing complex networks, as a key tool in our study. All methods were implemented using Python 3.8.18. For computing evaluation metrics, we utilized the sklearn.metrics module. Graph embeddings were facilitated by node2vec 0.4.6. We focus on comparing two different settings, i.e., a baseline using only graph embeddings and combining graph embeddings with time series contextual information. To ensure the reproducibility of our proposed settings, we have made the code available in a GitHub repository [24].
Section V-A shows detection results using a baseline model with only graph embeddings, demonstrating significant variation in performance with different $\omega$ and $\delta$ combinations using time-based windows (see Fig. 8). Section V-B evidences how incorporating time series features with graph embeddings enhances detection capabilities improving AUC-ROC values across various attacks (see Fig. 8) using time-based windows. Section V-C provides a comparative analysis of detection settings showing that combining graph embeddings with time series features generally enhances detection performance, supported by detailed AUC-ROC metrics across different attack scenarios (see Table II). Finally, Section V-D investigates the effect of using more granular windows based on the number of samples. We show that even with more granular window lengths, the proposed framework effectively detects masquerade attacks (see Table V and Figs. 10 and 10).
@@ V-A Baseline Model with Graph Embeddings Only
Our evaluation begins with the establishment of a baseline model that relies solely on graph embeddings, i.e., focusing only on the topology of the MSGs. This baseline model provides a fundamental comparison point and is essential for assessing the effectiveness of incorporating additional features. Fig. 8 shows the performance metrics for this baseline, serving as the benchmark for our subsequent enhancements. Specifically, the AUC-ROC values for ‘correlated signal attack’ reach an optimal performance of 0.98 with a $\omega$ of 8 seconds and an $\delta$ of 1 second, showcasing strong model effectiveness in this configuration. In contrast, ‘reverse light on attack’ achieves its highest AUC-ROC of 0.97 with $\omega=7$ seconds and $\delta=4$ seconds, respectively, demonstrating its dependence on extended time frames. In general for the remaining attacks, we notice that high classification results can be obtained for $\omega$ and $\delta$ of several seconds. These insights confirm that adjusting the configurations to suit different types of attacks is crucial for maximizing detection capabilities. The variance in performance across different settings shows the importance of fine-tuning the sliding window parameters to enhance the efficacy of anomaly detection systems in vehicular networks.
@@
Figure 7: AUC-ROC across different window size and offset combinations for masquerade attacks detected by the baseline model using only graph embeddings.
@@
Figure 8: AUC-ROC across different window size and offset combinations for masquerade attacks detected by incorporating both graph embeddings and time series features.
V-B Enhanced Model with Graph Embeddings and Time Series Features
@@
Building upon our baseline, we incorporate time series features of CAN signals alongside the graph embeddings. This setting harnesses both the structural patterns and the temporal dynamics of CAN frames aiming to achieve a more comprehensive anomaly detection. Fig. 8 depicts the performance metrics of this enhanced model. The inclusion of time series data significantly improves detection capabilities, as evidenced by the AUC-ROC scores. For ‘correlated signal attack’, the model achieves an AUC-ROC peak of 0.99 at a $\omega$ of 3 seconds and an $\delta$ of 3 seconds. This suggests a strong model response to quick changes in CAN signal patterns. Similarly, ‘reverse light on attack’ records its highest AUC-ROC at 0.99 for a $\omega$ of 7 seconds and an $\delta$ of 4 seconds, indicating effective monitoring of critical engine parameters. In general, we notice that when integrating time series features, the required time to get optimal performance is lower than when using only graph embeddings. Specifically, across the five masquerade attacks, the average optimal $\omega$ decreases from 9.8 seconds to 6 seconds, while the average optimal $\delta$ increases from 3.2 seconds to 4.8 seconds when combining graph embeddings with time series features. This 38.8% reduction in $\omega$ coupled with a 50% increase in $\delta$ allows for more granular analysis of CAN messages while spacing the analyses further apart, enabling our method to capture detailed patterns more effectively.
@@
TABLE II: Comparison of detection settings based on the AUC ROC heatmaps. The table shows the mean ($\mu$), standard deviation ($\sigma$), median ($\eta$), minimum ($\min$), and maximum ($\max$) of the AUC ROC for each setting across different attack scenarios, as well as the average ($\overline{\max}$) and standard deviation ($\sigma_{\max}$) of the maximum values obtained by each attack and experiment type, either embeddings only or embeddings with time series information.
Experiment Correlated signal
attack
Max engine coolant
temperature attack
Max speedometer
attack
Reverse light off
attack
Reverse light on
attack
Experiment max
Graph embeddings only $\mu=0.94$ $\sigma=0.02$ $\eta=0.93$ min = 0.90 (15,8) max = 0.98 (8,1) $\mu=0.91$ $\sigma=0.04$ $\eta=0.92$ min = 0.84 (15,14) max = 0.98 (9,8) $\mu=0.91$ $\sigma=0.03$ $\eta=0.91$ min = 0.85 (4,1) max = 0.98 (14,2) $\mu=0.93$ $\sigma=0.02$ $\eta=0.93$ min = 0.89 (6,3) max = 0.98 (11,1) $\mu=0.92$ $\sigma=0.03$ $\eta=0.92$ min = 0.88 (2,1) max = 0.97 (7,4) $\overline{\text{max}}=0.98$ $\sigma_{\text{max}}=0.00$
@@
Graph embeddings + Time series features $\mu=0.96$ $\sigma=0.02$ $\eta=0.96$ min = 0.92 (6,3) max = 0.99 (3,3) $\mu=0.94$ $\sigma=0.04$ $\eta=0.94$ min = 0.86 (8,1) max = 0.99 (7,4) $\mu=0.94$ $\sigma=0.03$ $\eta=0.93$ min = 0.87 (13,10) max = 0.99 (8,8) $\mu=0.96$ $\sigma=0.02$ $\eta=0.96$ min = 0.91 (6,3) max = 0.99 (5,5) $\mu=0.94$ $\sigma=0.03$ $\eta=0.94$ min = 0.90 (4,1) max = 0.99 (7,4) $\overline{\text{max}}=0.99$ $\sigma_{\text{max}}=0.00$

### v2

@@ IV-B Time Series Extraction and Analysis
Concurrently with the creation of the MSGs, raw CAN messages are decoded using a DBC as described in [cssElectronicsDBC]. In doing so, CAN data is transformed into a time series format to capture detailed patterns of the signals contained in every CAN ID per sliding window. We used CAN-D [Verma2021] to extract timeseries from CAN logs. Up to the time of this writing, CAN-D is still the state-of-the-art method for CAN reverse engineering [10092880]. Fig. 4 shows a decoded representation of the signals representing the four wheels’ speed of a vehicle [verma2022addressing]. The plot is generated by translating the CAN frames into a continuous time series format, which captures the network activity over time. This representation facilitates a detailed characterization of masquerade attacks within CANs highlighting variations and trends that may not be evident relying on raw CAN data.
Figure 4: Speed signals of four wheels encoded in node ID 1760. X-axis: Time (ms); Y-axis: Signal value.
@@ IV-C Graph Annotation
We enhance each MSG by annotating its nodes with statistical attributes derived from corresponding time series data. For each CAN ID, we compute the mean and standard deviation of its associated signals—features chosen for their simplicity and constant-time complexity, as demonstrated in prior work [shahriar_cantropy]. This annotation occurs alongside MSG construction, allowing the model to jointly capture structural and temporal dynamics. Fig. 5 illustrates the annotation process, and Algorithm 3 provides implementation details.
Algorithm 3 Node annotation with statistical attributes
0: CAN data $\mathcal{D}$, sliding windows $\mathcal{W}$, signal mappings $\mathcal{S}$
0: Annotated graphs $\{G_{t}\}$ with statistical node features
1:for each window $w\in\mathcal{W}$ do
2:  for each signal $s\in\mathcal{S}$ do
3:   Extract signal values $v_{s}$ from $\mathcal{D}$ within window $w$
4:   Interpolate missing values in $v_{s}$ using linear interpolation
5:   Compute mean $\mu_{s}$ and standard deviation $\sigma_{s}$ of $v_{s}$
6:   Annotate corresponding node in $G_{w}$ with $(\mu_{s},\sigma_{s})$
7:  end for
8:end for
9:return Annotated graphs $\{G_{t}\}$
@@ IV-D Graph Embeddings
To model the complex interactions in CANs, we leverage graph-based techniques, specifically graph embeddings, which map network structures into low-dimensional vectors for ML.
Graph embeddings map nodes, edges, or entire graphs into a lower-dimensional vector space, preserving their structural and relational information [grover2016node2vec, wu2020comprehensive]. In our study, we focus on shallow graph embeddings, specifically the node2vec algorithm [grover2016node2vec], which is well-suited for capturing the nuances of weighted and directed graphs characteristic of MSGs.
node2vec generates embeddings by simulating random walks on the graph, controlled by two key parameters: the return parameter $p$ (likelihood of revisiting a node) and the in-out parameter $q$ (preference for exploring local vs. distant neighborhoods). These parameters allow the algorithm to balance the exploration of immediate connectivity and broader network contexts, making it particularly effective for modeling CAN bus data [wei2024robustnessgraphembeddingmethods].
We configure node2vec with 64 dimensions ($d$) to balance computational efficiency and feature richness [Gu2021], a walk length ($l$) of 15 to capture multi-hop relationships, and 100 walks per node ($r$) to ensure thorough exploration of the network. Additionally, we set the return parameter ($p$) to 1.5 and the in-out parameter ($q$) to 0.5, emphasizing local ECU interactions while retaining global CAN topology. This configuration ensures that the embeddings capture both the immediate communication patterns between ECUs and the broader structure of the CAN.
Unlike prior works that focus solely on node embeddings [Islam2020, Jedh2021], we construct whole-graph embeddings by averaging node embeddings within sliding windows. This aggregation process provides a unified representation of the network’s behavior, capturing both structural and temporal patterns of CAN frames. Specifically, we stack the embeddings per node in the MSGs, labeling each graph to indicate normal or anomalous activity in the CAN. This approach allows us to frame our intrusion detection problem as a graph classification task [Zhong2024], combining both structural and temporal information for enhanced detection capabilities.
The relatively low embedding dimension (64) has been shown to be sufficient for capturing rich graph-related properties while maintaining computational efficiency [Gu2021], which is critical for near real-time detection of masquerade attacks in CANs. By enriching the embeddings with time-series features (e.g., mean, standard deviation) from decoded signals, we ensure a context-aware analysis of masquerade attacks, further enhancing the robustness of our framework.
@@ IV-E Supervised Learning
To evaluate the effectiveness of our feature representation and overall anomaly detection framework, we employ supervised learning techniques. We experiment with two ensemble learning methods: Random Forest (RF) [breiman2001random] and XGBoost [10.1145/2939672.2939785]. These decision tree-based ensemble learning methods are renowned for their high accuracy, scalability, and robustness, making them exceptionally suitable for dealing with the high-dimensional space characteristic of CAN data. The RF model is instantiated with 100 trees, and depth is capped at 20 to prevent overfitting. A combination of a minimum samples split of 6 and minimum samples leaf of 2 is used to control the growth of the trees further. As the training dataset exhibits a 1:10 ratio between the minority (attack) and majority (normal) classes, we balance the dataset using the synthetic minority over-sampling technique (SMOTE) [10.5555/1622407.1622416]. A 5-fold stratified k-fold cross-validation is performed to ensure the model’s generalizability during training. The XGBoost model yielded similar results to the RF classifier in our experiments. Despite the comparable performance, we opted to report only results for the RF classifier due to its simpler interpretability and less intensive hyperparameter tuning process, which aligns well with the goals of this study.
@@ IV-F Evaluation Setup
Evaluation is performed at the sliding window level. In doing so, we verify the proportion of sliding windows coinciding with an attack region in the testing captures, which refers to the recorded CAN bus data used for testing. We found that for the combinations of $(\omega_{t},\delta_{t})$ or $(\omega_{s},\delta_{s})$, the proportion of windows containing attacks tended to be balanced, i.e., greater than 40%. To quantitatively assess the performance of our anomaly detection framework, we use the area under the receiver operating characteristic curve (AUC-ROC) [FAWCETT2006861]. The AUC-ROC is a comprehensive evaluation metric for classification problems across various threshold settings. It reflects the model’s ability to distinguish between classes, with higher values indicating better performance. The ROC curve is a plot of the true positive rate (TPR) against the false positive rate (FPR) at different threshold levels. The TPR, also known as sensitivity or recall, measures the proportion of actual positives correctly identified, while the FPR measures the proportion of actual negatives incorrectly identified as positive. Mathematically, the AUC-ROC is defined as: $\text{AUC-ROC}=\int_{0}^{1}\text{TPR}(\text{FPR}^{-1}(x))\,dx.$
In practical terms, the AUC-ROC can be interpreted as the probability that the classifier will rank a randomly chosen positive instance higher than a randomly chosen negative instance. This interpretation is particularly useful in our context, where we aim to detect subtle masquerade attacks embedded within normal CAN traffic.
Performance Metrics: We evaluated computational performance using the testing time per window (TTW) metric by Nichelini et al. [nichelini2023canova], which quantifies the amount of time required to process each window across attack scenarios. We define TTW as:
$\text{TTW}=\frac{\text{Total Detection Time}}{\text{Number of Windows}}$
TTW provides a practical means of evaluating the computational feasibility of our detection framework in real-time settings. All experiments were conducted on a 12th Gen Intel Core i9-12900HK CPU with an NVIDIA RTX 3080 Ti GPU and 32GB RAM.
@@ IV-G Dataset
Our framework is evaluated on the Real ORNL Automotive Dynamometer (ROAD) dataset [verma2022addressing], collected from a physical vehicle at Oak Ridge National Laboratory (ORNL). This dataset includes physically verified fabrication and simulated masquerade attacks, offering a realistic benchmark for CAN security methods. It is distinguished by its high-quality CAN data and credible attack scenarios across diverse driving conditions. The dataset comprises 3.5 hours of recorded data—3 hours for training and 30 minutes for testing. The test data incorporates five masquerade attacks including correlated signal, max engine coolant, max speedometer, reverse light off/on attacks. These attacks are designed to manipulate specific vehicle states. This dataset enables us to conduct a comprehensive evaluation of our framework under a variety of attack conditions. Table II summarizes the durations, injection intervals, and impacts of the masquerade attacks.
TABLE II: COMPARISON OF RELATED WORKS ON INTRUSION DETECTION IN CANS.
Attack Name Duration (s) Injection Interval (s) Description correlated_signal_1 33.10 [9.19, 30.05] Injects varying values for wheel speeds, causing the vehicle to halt. correlated_signal_2 28.23 [6.83, 28.23] correlated_signal_3 16.96 [4.32, 16.96] max_speedometer_1 88.02 [42.01, 66.45] Injects the maximum value to be displayed on the speedometer. max_speedometer_2 59.70 [16.01, 47.41] max_speedometer_3 86.77 [9.52, 70.59] max_engine_coolant 25.88 [19.98, 24.17] Injects the maximum value, triggering the coolant warning light. reverse_light_off_1 28.11 [16.63, 23.35] Toggles the reverse light irrespective of the actual gear position. reverse_light_off_2 40.67 [13.17, 36.88] reverse_light_off_3 57.88 [16.52, 40.86] reverse_light_on_1 54.85 [18.93, 38.84] Toggles the reverse light irrespective of the actual gear position. reverse_light_on_2 72.02 [20.41, 57.30] reverse_light_on_3 64.26 [23.07, 46.58]
@@ V Results
We evaluate our framework on the ROAD dataset [verma2022addressing] under two settings: (1) a baseline using graph embeddings only and (2) an enhanced configuration combining graph embeddings with time series features. All methods were implemented in Python 3.8.18, using NetworkX for graph operations, sklearn.metrics for evaluation, node2vec 0.4.6 for embeddings, and timeit for computing TTW. To support reproducibility, our code is publicly available at [GraphML-CONTROLLER-AREA-NETWORK-Attack-Detection].
Section V-A shows detection results using a baseline model with graph embeddings only, demonstrating significant variation in performance across different $\omega_{t}$ and $\delta_{t}$ combinations using time-based windows (see Fig. 7). Section V-B evidences how incorporating time series features with graph embeddings enhances detection capabilities, improving AUC-ROC values across various attacks (see Fig. 7). Section V-C provides a comparative analysis of detection settings (see Table III) and also presents TTW as a proxy for computational overhead across attack types (see Table V). This analysis supports the feasibility of our approach for time-sensitive applications and highlights trade-offs introduced by different window configurations. Section V-D presents a comparison with state-of-the-art (SOTA) methods, contextualizing our method’s performance against existing approaches (see Table VIII). Finally, Section V-E investigates the effect of using more granular windows based on the number of samples. We show that even under these fine-grained configurations, the framework maintains strong detection performance (see Table IX).
@@ V-A Baseline Model with Graph Embeddings Only
Our evaluation begins with the establishment of a baseline model that relies solely on graph embeddings, i.e., focusing only on the topology of the MSGs. This baseline model provides a fundamental comparison point and is essential for assessing the effectiveness of incorporating additional features. Fig. 7 shows the performance metrics for this baseline serving as the benchmark for our subsequent enhancements. Specifically, the AUC-ROC values for correlated_signal_attack reach an optimal performance of 0.98 with a $\omega_{t}$ of 8 seconds and an $\delta_{t}$ of 1 second, showcasing strong model effectiveness in this configuration. In contrast, reverse_light_on_attack achieves its highest AUC-ROC of 0.97 with $\omega_{t}=7$ seconds and $\delta_{t}=4$ seconds, respectively, demonstrating its dependence on extended time frames. In general for the remaining attacks, we notice that high classification results can be obtained for $\omega_{t}$ and $\delta_{t}$ of several seconds. These insights confirm that adjusting the configurations to suit different types of attacks is crucial for maximizing detection capabilities. The variance in performance across different settings shows the importance of fine-tuning the sliding window parameters to enhance the efficacy of anomaly detection systems in vehicular networks.
@@
Figure 6: AUC-ROC across different $\omega_{t}$ and $\delta_{t}$ combinations for masquerade attacks detected by the baseline model using only graph embeddings.
@@
Figure 7: AUC-ROC across different $\omega_{t}$ and $\delta_{t}$ combinations for masquerade attacks detected by incorporating both graph embeddings and time series features.
V-B Enhanced Model with Graph Embeddings and Time Series
@@
Building upon our baseline, we incorporate time series features of CAN signals alongside the graph embeddings. This setting harnesses both the structural patterns and the temporal dynamics of CAN frames aiming to achieve a more comprehensive anomaly detection. Fig. 7 depicts the performance metrics of this enhanced model. The inclusion of time series data significantly improves detection capabilities, as evidenced by the AUC-ROC scores. For correlated_signal_attack, the model achieves an AUC-ROC peak of 0.99 at a $\omega_{t}$ of 3 seconds and an $\delta_{t}$ of 3 seconds. This suggests a strong model response to quick changes in CAN signal patterns. Similarly, reverse_light_on_attack records its highest AUC-ROC at 0.99 for a $\omega_{t}$ of 7 seconds and an $\delta_{t}$ of 4 seconds, indicating effective monitoring of critical engine parameters. In general, we notice that when integrating time series features, the required time to get optimal performance is lower than when using only graph embeddings. Specifically, across the five masquerade attacks, the average optimal $\omega_{t}$ decreases from 9.8 seconds to 6 seconds, while the average optimal $\delta_{t}$ increases from 3.2 seconds to 4.8 seconds when combining graph embeddings with time series features.
Offset Interpretation and Design Rationale: The offset value is not equivalent to a detection delay—rather, it is a mechanism to ensure continuous attack coverage by allowing consecutive windows to overlap. Without this, attacks occurring near the window boundaries could be missed. Increasing the offset (e.g., to 6s) risks skipping attacks entirely, while reducing it (e.g., to 1s) significantly increases computational overhead. Our configuration optimally balances attack coverage and processing feasibility while ensuring the IDS remains responsive. This 38.8% reduction in $\omega_{t}$ (from 9.8 s to 6 s) coupled with a 50% increase in $\delta_{t}$ (from 3.2 s to 4.8 s) allows for more granular analysis of CAN messages while spacing the analyses further apart, enabling our method to capture detailed patterns more effectively.
@@
TABLE III: COMPARISON OF DETECTION SETTINGS BASED ON AUC-ROC. THE TABLE REPORTS THE MEAN ($\mu$), STANDARD DEVIATION ($\sigma$), MEDIAN ($\eta$), MINIMUM (min), AND MAXIMUM (max) AUC-ROC VALUES FOR EACH ATTACK SCENARIO. IT ALSO INCLUDES THE AVERAGE ($\overline{\text{max}}$) AND STANDARD DEVIATION ($\sigma_{\text{max}}$) OF THE MAXIMUM VALUES OBTAINED ACROSS ALL EXPERIMENT TYPES—GRAPH EMBEDDINGS ONLY, TIME SERIES ONLY, AND THE HYBRID APPROACH.
Experiment Correlated signal Max speedo- meter Max engine coolant Reverse light off Reverse light on Experiment max
Graph embeddings only $\mu=0.94$ $\sigma=0.02$ $\eta=0.93$ min = 0.90 (15,8) max = 0.98 (8,1) $\mu=0.91$ $\sigma=0.03$ $\eta=0.91$ min = 0.85 (4,1) max = 0.98 (14,2) $\mu=0.91$ $\sigma=0.04$ $\eta=0.92$ min = 0.84 (15,14) max = 0.98 (9,8) $\mu=0.93$ $\sigma=0.02$ $\eta=0.93$ min = 0.89 (6,3) max = 0.98 (11,1) $\mu=0.92$ $\sigma=0.03$ $\eta=0.92$ min = 0.88 (2,1) max = 0.97 (7,4) $\overline{\text{max}}=0.98$ $\sigma_{\text{max}}=0.00$
Time series features only $\mu=0.82$ $\sigma=0.05$ $\eta=0.83$ min = 0.76 (13,7) max = 0.90 (10,4) $\mu=0.85$ $\sigma=0.04$ $\eta=0.86$ min = 0.80 (9,6) max = 0.91 (11,5) $\mu=0.79$ $\sigma=0.06$ $\eta=0.80$ min = 0.72 (14,9) max = 0.89 (12,5) $\mu=0.81$ $\sigma=0.05$ $\eta=0.83$ min = 0.75 (13,8) max = 0.88 (9,4) $\mu=0.84$ $\sigma=0.03$ $\eta=0.85$ min = 0.79 (8,5) max = 0.90 (10,5) $\overline{\text{max}}=0.90$ $\sigma_{\text{max}}=0.01$
@@
Graph embeddings + time series features $\mu=0.96$ $\sigma=0.02$ $\eta=0.96$ min = 0.92 (6,3) max = 0.99 (3,3) $\mu=0.94$ $\sigma=0.03$ $\eta=0.93$ min = 0.87 (13,10) max = 0.99 (8,8) $\mu=0.94$ $\sigma=0.04$ $\eta=0.94$ min = 0.86 (8,1) max = 0.99 (7,4) $\mu=0.96$ $\sigma=0.02$ $\eta=0.96$ min = 0.91 (6,3) max = 0.99 (5,5) $\mu=0.94$ $\sigma=0.03$ $\eta=0.94$ min = 0.90 (4,1) max = 0.99 (7,4) $\overline{\text{max}}=0.99$ $\sigma_{\text{max}}=0.00$